a train leves at 9 am going 60 mph and another train leaves at 12 noon going 75 mph. how long until the second train catches the first

Found 2 solutions by mananth, Alan3354:
Answer by mananth(16949)

train A leaves 9.00am at 60 mph.
Train B leaves at 12 noon at 75 mph
..
Before train B leaves train A has traveled 60*3 = 180 miles.
From this point let train A travel x miles before they meet.
..
So train B travels x+180 miles to catch up.
..
time taken by train A = x/60
time taken by train B = x+180 / 75
..
x/60 = x+180 / 75
75x = 60(x+180)
75x=60x +10800
15x=10800
x= 720 miles.
Train B has to travel 720 miles to catch up.
its speed is 75 mph
time taken = 720/75
=9.6 hours

Answer by Alan3354(69443)

a train leves at 9 am going 60 mph and another train leaves at 12 noon going 75 mph. how long until the second train catches the first
------------------------
In 3 hours, the 1st train is 180 miles away (3*60)
The 2nd train gains on it at 15 mph (75 - 60)
180/15 = 12 hours